An ox and a mule were comparing the number of sacks they had to carry. The ox said, "right now i have three times as many sacks as you have. If i gave you four of my sacks i would still have twice as many as you would have then". How many sacks is each carrying?
Answer by richwmiller(17219)

x=3m
x-4=2*(m+4)
3m-4=2m+8
m=12
x=36
check
36-4=2*(12+4)
32=32
ok